ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 90, 1995

Aharoni, R. and M. Loebl Strongly perfect infinite graphs  81

Alcalde Cuesta, F. and G. Hector Intégration symplectique des variétés de Poisson régulières

125

Asada M. and H. Nakamura On graded quotient modules of mapping class groups of surfaces

93

Bartoszek, W. On convolution powers on semidirect products  115

Baston, V. and A. Beck Generalizations in the linear search problem

301

Beck, A. See V. Baston

Bilu, Y. Effective analysis of integral points on algebraic curves  235
Bogdanović, S. and M. Ćirić Orthogonal sums of semigroups  423

Ćirić, M. See S. Bogdanović

Erdős, P., A. Sárközy and V. T. Sós On sum sets of Sidon sets, II 221

Faltings, G. and B. W. Jordan Crystalline cohomology and

GL(2,ℚ) 1

Grabarnik, G. and A. Katz Ergodic type theorems for finite von Neumann algebras

403

Harboure, E., R. A. Macías, C. Segovia and J. L. Torrea Some estimates for maximal functions on Köthe function spaces

349

Hector, G. See F. Alcalde Cuesta

Hernández, F. L. and B. Rodríguez-Salinas Lattice-embedding L^{p} into Orlicz spaces

167

Jordan, B. W. See G. Faltings

Katz, A. See G. Grabarnik

Krattenthaler, C. On bideterminantal formulas for characters of classical groups

67

Loebl, M. See R. Aharoni

Macías, R. A. See E. Harboure

Mozes, S. Actions of Cartan subgroups  253

Nakamura, H. See M. Asada Nogueira, A. The three-dimensional Poincaré continued fraction

algorithm

373

Rodríguez-Salinas, B. See F. L. Hernández

Sárközy, A. See P. Erdős

Sastry, P. and A. Yekutieli On residue complexes, dualizing sheaves and local cohomology modules

325

Schmidt, K. Invariant measures for certain expansive ℤ^{2}-actions  295

Segovia, C. See E. Harboure

Shustin, E. Super-high-frequency oscillations in a discontinuous dynamic system with time delay

199

Song, Y. On the relative Wall–Witt groups  189

Sós, V. T. See P. Erdős

Torrea, J. L. See E. Harboure

Yekutieli, A. See P. Sastry